Problem4:

25 Ω b 1 = 0, T 80 Ω a 100 Ω + + + 100 V 240 Ω υο 80 μF 275 V Υ 16 mH

(b) Find Il(s). Find i(t). (c) Find the zero-input and the zero-state components of iz(t) for t > 0.
Answer #1

Eko son) louh 11 (1) as VIEI 100 zuo? suur il(t) 16m H at steady state 100 275 luo 2401 Vco) EL(0-) apply kch at u 100-v- V V2 5 100 (727513 3 colo mit Ls Jol 20 12.5x103 Vas I 1275 1.25L Inmitt 0.0L 0.0165 2. (20+ + 12.5X103 +0.0165) I (1) + 150 + sTisl= (6250 +5 ) 0:3 (3 462532 4/6258 5625 808 2( )= 0.8 (+625) + (3+625) 4(658 (32 +62553 € (bers? It 625 7.2 x 625 IL S 0.8